# COSMIC: Mutual Information for Task-Agnostic Summarization Evaluation

### v2

@@ Abstract
Assessing the quality of summarizers poses significant challenges. In response, we propose a novel task-oriented evaluation approach that assesses summarizers based on their capacity to produce summaries that are useful for downstream tasks, while preserving task outcomes. We theoretically establish a direct relationship between the resulting error probability of these tasks and the mutual information between source texts and generated summaries. We introduce COSMIC as a practical implementation of this metric, demonstrating its strong correlation with human judgment-based metrics and its effectiveness in predicting downstream task performance. Comparative analyses against established metrics like BERTScore and ROUGE highlight the competitive performance of COSMIC.
@@
[…]
@@ 1 Introduction
Standard automatic evaluation methods for summarization rely on the idea that a good summary should have some (semantic) overlap with either a gold standard or the source text El-Kassas et al. (2021); Allahyari et al. (2017). They leverage similarity metrics such as BLEU Papineni et al. (2002) or ROUGE Lin (2004) to evaluate the quality of the generated summaries. However, these methods often do not correlate well with human judgements Kryscinski et al. (2020); Kocmi et al. (2021).
To enhance alignment with human judgment and capitalize on recent advancements in large language models, recent efforts have concentrated on learned metrics for scoring summaries Zhang et al. (2020b); Rei et al. (2020); Liu et al. (2023). For instance, Clark et al. (2023) introduced six new learnt metrics by finetuning a pretrained MT5 model Xue et al. (2021) to predict human judgement along different axes.
@@
In this work, we introduce two modifications to the aforementioned paradigm. First, instead of evaluating the quality of each output summary individually, we propose to evaluate the probability distribution of the summary induced by a summarizer. Second, we adopt a task-oriented evaluation setup where summaries are meant to allow an agent to perform downstream tasks without reading the longer source text. For example, if a political advisor drafts a briefing on a subject to enable a politician to make informed decisions, our evaluation considers the advisor successful if the decisions made using the summary align with those made using the initial source text Pu et al. (2023); Vanderwende et al. (2007).
@@
Many summarization techniques typically operate under the assumption that the downstream task is unknown during the summarization process, yet an implicit notion of summary quality exists. However, this assumption is seldom explicitly addressed or substantiated. In our task-oriented approach, we provide an information-theoretic rationale for the existence of such a metric. We demonstrate that it essentially involves assessing the mutual information (MI) between the distribution of the source texts and the distribution of the summaries generated by a given summarizer.
@@
We evaluate the quality of our approach in two ways. We show that summarizers that induce a summary distribution with higher MI with the source texts’ distribution are higher quality in the following sense — they tend to produce summaries that preserve outcomes on downstream tasks as compared to using the source texts. Additionally, we compare the MI to metrics trained on human judgements and show that it displays consistent correlations.
Our results are consistent with and extend previous work that leverages MI to understand relationships in datasets Ethayarajh et al. (2022); Bugliarello et al. (2020), predictive performance of representations Sui et al. (2023) and tool to construct or evaluate representations Kim et al. (2022), models and generative processes.
@@ 2 Related Work
Assessing the quality of summaries poses a unique challenge due to the subjective nature of ‘quality’, influenced by factors such as the audience, topic, and intended purpose of the summary. Even expertly crafted human summaries considered "gold standard" in one context may be perceived as subpar in a different setting Saziyabegum and Sajja (2017); Indu and Kavitha (2016).
Reference-free summary evaluation. Reference-free evaluation methods mostly rely on comparing the content of the summary with the content of the source text Louis and Nenkova (2013); El-Kassas et al. (2021) and they rely on common overlap metrics such as ROUGE Lin (2004), BLEU Papineni et al. (2002) or BERTScore Zhang et al. (2020b). However, most reference-free metrics show some important limitations Deutsch et al. (2022): they lack theoretical grounding. We argue that our proposed method avoids these shortcomings as it offers a solid theoretical framework to understand summarization evaluation.
@@ 3.2 Background: Information Theory
Information theory Cover and Thomas (2006) provides several tools for analyzing data and their associated probability distributions, including entropy and MI. These metrics are typically defined based on a "true" probability distribution, denoted as $p(c)$, or the joint probability density function $p(\mathbf{t},\mathbf{s})$ which may not be known but governs the behaviour of random variables $\mathbf{T}$ and $\mathbf{S}$. The fundamental concept in information theory is surprisal, defined as $H(C=c)=-\log p(c)$, and its expected value is termed entropy:
@@
3.3 Background: Classification-based evaluation
Most summarization methods operate under the assumption that the downstream task of interest is unknown during the generation process, relying instead on a generic notion of summary quality. We assume in this work the goal of being able to perform similarly in terms of classification error on both the original texts $\mathbf{T}$ and the resulting summaries $\mathbf{S}$. Next, we formalize this evaluation metric based on the assumption of an unknown downstream task.
Let $c:\Omega^{*}\rightarrow\{1,\dots,m\}$ denote the target concept of interest which can be extracted from the initial texts $\mathbf{T}$ by applying $C\triangleq c(\mathbf{T})$; and similarly, let $\widehat{c}(\mathbf{S})$ denote the predicted concept from the summaries $\mathbf{S}$ according to the underlying model $p_{\theta}(\mathbf{s}|\mathbf{t})$. The evaluation of the summarizer’s quality with respect to the downstream task (unknown from the summarization model) can be assessed using the error probability, as illustrated in Figure 1. In other words, it involves the classifier, determining the average probability of the extracted concept $\widehat{c}(\mathbf{S})$ differing from the original concept $c(\mathbf{T})$ in the source text:
@@ 3.3 Background: Classification-based evaluation
where the expectation is taken with respect to the joint distribution of text and summary $(\mathbf{T},\mathbf{S})$ based on the underlying summarization model. Interestingly, it is not difficult to check by data-processing inequality that (4): $I(C;\mathbf{S})\leqslant I(C;\mathbf{T})$ and thus, the output summaries $\mathbf{S}$ by the summarization model may not preserve all relevant information necessary to predict $C$ from $\mathbf{S}$, unless the summarization model is aware of the downstream task. However, in practice, identifying the set of relevant downstream tasks for various texts is challenging. Consequently, (6) does not provide a satisfactory evaluation metric for summarization systems. This observation raises the central question studied in this work: How much information about any downstream task is preserved in the summaries?
4 A Quality Measure for Task-Agnostic Summarization
@@ 4.1 Theoretical results
In this section, we rigorously motivate the evaluation of summarization systems by measuring the MI between texts and the resulting summaries. Let’s assume the existence of a random variable $\mathbf{T}$ representing source texts, which follows an unknown distribution inherent to the source text domain. Given a stochastic summarization system $p_{\theta}(\mathbf{s}|\mathbf{t})$ as defined in expression (1), we denote $\mathbf{S}\sim p_{\theta}(\mathbf{s}|\mathbf{T})$ the random variable representing summaries generated by the summarization system for these sources texts $\mathbf{T}$.
Consider the task $c(\mathbf{T})$ that we intend to execute on the source texts, where $C=c(\mathbf{T})$ represents the random variable denoting the outcomes of this task on source texts. We assume that the same task can be performed on the summaries and let $\widehat{c}(\mathbf{S})$ denote the corresponding prediction. Intuitively, if we can accurately predict the outcome of $C$ from the summaries, then we will say that the summarization system is high-quality since it preserves the necessary information for performing the task. The next proposition frames our information-theoretic bounds on the performance of any arbitrary downstream task. In particular, it shows that the error probability $P_{e}(c,\widehat{c},\theta)$ as defined in (6) can be upper and lower bounded by the MI between the texts and the summaries.
@@
Let $C=c(\mathbf{T})$ denote the underlying concept variable and let $\widehat{c}(\mathbf{S})$ be the Bayes predictor of $C$ observing the output summaries $\mathbf{S}$, based on the underlying summarization model $p_{\theta}(\mathbf{s}|\mathbf{t})$. The error probability satisfies:
@@
The upper bound (7) relies on the fact that the predictor $\widehat{c}(\mathbf{S})$ is the optimal (Bayes) classifier for which the error probability admits a well-known expression. The lower bound (8) uses data-processing inequality which implies that $I(\mathbf{T};\mathbf{S})\geqslant I\big{(}c(\mathbf{T});\widehat{c}(\mathbf{S})\big{)}$ and the definition of the rate-distortion function evaluated in the loss $\ell_{01}(c,\widehat{c})$ noticing that its expectation yields the error probability $P_{e}(c,\widehat{c},\theta)$ as defined in (6). For further details, see Appendix A. ∎
@@
The bounds in Proposition 1 imply that the error probability of any task predicted by observing the summaries is lower- and upper-bounded by functions of the MI between the text and the summaries $I(\mathbf{T};\mathbf{S})$. More precisely, the upper bound (7) is a monotonically decreasing function of the MI while the lower bound (8) is a non-increasing function in the MI. The lower bound can be further simplified by evaluating the rate-distortion function, as shown in Appendix B. In other words, greater MI corresponds to improved prediction performance on the summaries. Conversely, when MI is limited, the error probability is bounded from above. Interestingly, the arguments in the bounds do not depend on the considered task, suggesting that the MI can be used as a task-agnostic metric to evaluate the quality of summaries.
In the next section, we propose a practical method to estimate MI. In Section 6, we empirically show that it correlates well with the performance of the downstream tasks with other human judgement-based metrics, and compare it to standard metrics. Discussions are relegated to Section 7.
@@ 4.2 Estimating MI from samples
Mutual information estimator. We rely on the KNIFE estimator Pichler et al. (2022) to estimate the differential entropy of the embeddings and then mutual information. This estimator effectively relies on Gaussian Mixtures with $K$ modes to fit the density functions and induces a soft-clustering for text generation evaluation Pillutla et al. (2021); Pimentel et al. (2023). We found experimentally that the number of modes $K$ did not impact the performance significantly. We report the results for $K=4$ (see Appendix D.1 for further details).
Embeddings. In order to obtain continuous representations of the source texts and the summaries, we mainly rely on the AnglE-embedders Li and Li (2023) as they are at the top of the MTEB leaderboard Muennighoff et al. (2023) with a rather small model. In addition, we experimented with different embedders from the sentence transformers library Reimers and Gurevych (2019), mainly the paraphrase and sentence similarity embeddings. We find that our method was robust to the choice of the embedder and thus report the results for the AnglE-embedders.
@@ 5 Experimental Settings
We present two evaluation frameworks. Initially, we show that MI effectively predicts whether a downstream task yields the same outcome when applied to the summaries and the source text. This validates our task-oriented evaluation approach for summaries. Furthermore, we contrast our metric with metrics trained to emulate human judgment across various dimensions. Remarkably, we observe that even without any training, our metric closely aligns with inherent human preferences.
Datasets. We select three well-known summarization datasets for the English language: CNN/DailyMail See et al. (2017); Hermann et al. (2015),XSum Narayan et al. (2018) and MultiNews Fabbri et al. (2019) and perform all evaluations on all datasets. We provide additional experiments in French and Spanish using the MLSUM Scialom et al. (2020) and XLSUM Hasan et al. (2021) datasets. We report mixed results due to the lack of efficient multilingual embedders in Section E.2.
Models. We evaluate numerous summarizers from the HuggingFace hub, relying on different backbones, pretraining methods and finetuned on different datasets. We conduct most of our experiments on the PEGASUS suite of models Zhang et al. (2020a), on BERT large models Lewis et al. (2019) and on DistilBERT models Shleifer and Rush (2020). In addition, we evaluated the models presented in the SEAHORSE benchmark Clark et al. (2023) through the generated summaries they proposed, which include MT5 models Xue et al. (2021), different variants of T5 models Raffel et al. (2023), and of the PaLM models Chowdhery et al. (2022). The reader is referred to Appendix C for the details on the models.

### v3

@@ Abstract
Assessing the quality of summarizers poses significant challenges—gold summaries are hard to obtain and their suitability depends on the use context of the summarization system. Who is the user of the system, and what do they intend to do with the summary? In response, we propose a novel task-oriented evaluation approach that assesses summarizers based on their capacity to produce summaries while preserving task outcomes. We theoretically establish both a lower and upper bound on the expected error rate of these tasks, which depends on the mutual information between source texts and generated summaries. We introduce COSMIC, a practical implementation of this metric, and demonstrate its strong correlation with human judgment-based metrics, as well as its effectiveness in predicting downstream task performance. Comparative analyses against established metrics like BERTScore and ROUGE highlight the competitive performance of COSMIC.
@@
Maxime Darrin1,2,3,4  Philippe Formont1,2,4,5 Jackie Chi Kit Cheung2,3,7Pablo Piantanida1,2,4,6 1International Laboratory on Learning Systems, 2Mila - Quebec AI Institute 3McGill University 4Université Paris-Saclay, 5École de technologie supérieure (ETS) 6CNRS, CentraleSupélec 7Canada CIFAR AI Chair maxime.darrin@mila.quebecphilippe.formont@mila.quebec jackie.cheung@mcgill.capablo.piantanida@mila.quebec
@@ 1 Introduction
Standard automatic evaluation methods for summarization rely on the idea that a good summary should have some (semantic) overlap with either a gold standard or the source text El-Kassas et al. (2021); Allahyari et al. (2017). They leverage similarity metrics such as BLEU Papineni et al. (2002) or ROUGE Lin (2004) to evaluate the quality of the generated summaries. However, these methods often do not correlate well with human judgements Kryscinski et al. (2020); Kocmi et al. (2021). To enhance alignment with human judgment and capitalize on recent advancements in large language models, recent efforts have concentrated on learned metrics for scoring summaries Zhang et al. (2020b); Rei et al. (2020); Liu et al. (2023). For instance, Clark et al. (2023) introduced six new learned metrics by finetuning a pretrained MT5 model Xue et al. (2021) to predict human judgment along different axes. Another line of research focuses on reconstruction-based metrics, such as BLANC Vasilyev et al. (2020) and Information Difference Egan et al. (2021), the former evaluating the actual reconstruction error of the source text with and without a summary, and the latter evaluating the information gain obtained when conditioning the generative model on the summary.
Standard evaluation methods, therefore, suffer from two major shortcomings. They rarely formally define a clear notion of quality and the methods used to evaluate said notion lack theoretical foundations. This leads to potential discrepancies between the intended evaluation and what is actually measured, leading to a lack of validity of the evaluation methods.
@@
We introduce a task-oriented evaluation setup where summaries are meant to allow an agent to perform downstream tasks without reading the longer source text. For example, if a political advisor drafts a briefing on a subject to enable a politician to make informed decisions, our evaluation considers the advisor successful if the decisions made using the summary align with those made using the initial source text Pu et al. (2023); Vanderwende et al. (2007). Providing such a well-defined notion of success enables careful analysis of the validity of the evaluation for different possible use cases. In addition, it enables us to perform mechanical evaluations of our method, reducing the noise from the evaluation process.
@@
Furthermore, many summarization techniques typically operate under the assumption that the downstream task is unknown during the summarization process, yet an implicit notion of summary quality exists. However, this assumption is seldom explicitly addressed or substantiated. In our task-oriented approach, we provide an information-theoretic rationale for the existence of such a metric. We demonstrate that it essentially involves assessing the mutual information (MI) between the source texts’ distribution and the summaries generated by a given summarizer.
This paper does not aim to study the syntactic, semantic, and pragmatic aspects of features of summary information, which are essential for capturing the rich notion of information in human communication. For instance, we do not account for the summaries’ fluency, grammatical correctness, or coherence. Instead, our approach emphasizes how effectively, in the best-case scenario, the output of a summarizer can be used to perform downstream tasks. The results in this paper suggest that the formal definition of mutual information successfully achieves this goal. Our approach leverages embeddings to abstract the surface form of text, following recent studies by Pillutla et al. (2021) and Pimentel et al. (2023).
@@
We evaluate the quality of our approach in two ways. First, we show that summarizers that induce a summary distribution with higher MI with the source texts’ distribution are higher quality in the following sense — they tend to produce summaries that preserve outcomes on downstream tasks as compared to using the source texts. Second, we compare the MI to metrics trained on human judgments and show that it displays consistent correlations. Our results are consistent with and extend previous work that leverages MI to understand relationships in datasets Ethayarajh et al. (2022); Bugliarello et al. (2020), predictive performance of representations Sui et al. (2023) and tool to construct or evaluate representations Kim et al. (2022), models and generative processes.
@@ 2 Related Work
Assessing the quality of summaries poses a unique challenge due to the contextual nature of ‘quality’, influenced by factors such as the audience, topic, and intended purpose of the summary. Even expertly crafted human summaries considered "gold standard" in one context may be perceived as subpar in a different setting Saziyabegum and Sajja (2017); Indu and Kavitha (2016).
Reference-free summary evaluation. Reference-free evaluation methods mostly rely on comparing the content of the summary with the content of the source text Louis and Nenkova (2013); El-Kassas et al. (2021) and they rely on common overlap metrics such as ROUGE Lin (2004), BLEU Papineni et al. (2002) or BERTScore Zhang et al. (2020b). However, most reference-free metrics show some important limitations Deutsch et al. (2022): they lack theoretical grounding. Work such as BLANC Vasilyev et al. (2020) and Information Difference Egan et al. (2021) are closely related to our work. The former evaluates the actual reconstruction error of the source text with and without a summary, while the latter evaluates the information difference evaluated by the generative model. However, both lack the theoretical justification we introduce in Section 3 and only evaluate the information based on the generative model. While these methods appear intuitive, they lack of theoretical support for their approaches, relying solely on empirical results and correlations with human evaluations. Thus, the results are often tied to a dataset and are affected by variance from the human evaluation. Conversely, we offer a theoretical framework applicable to any dataset and measure a tangible, well-defined success criterion: do the tasks yield the same output when performed on summaries as they do on the source texts?
@@ 3.2 Background: Information Theory
Information theory Cover and Thomas (2006) provides several tools for analyzing data and their associated probability distributions, including entropy and MI. These metrics are typically defined based on a "true" probability distribution, denoted as $p(c)$, or the joint probability density function $p(\mathbf{t},\mathbf{s})$ which may not be known but governs the behavior of random variables $\mathbf{T}$ and $\mathbf{S}$. The fundamental concept in information theory is surprisal, defined as $H(C=c)=-\log p(c)$, and its expected value is termed entropy:
@@
3.3 A task-oriented evaluation setting
Most summarization methods operate under the assumption that the downstream task of interest is unknown during the generation process, relying instead on a generic notion of summary quality. We assume in this work that the goal is to be able to perform similarly in terms of classification error on both the original texts $\mathbf{T}$ and the resulting summaries $\mathbf{S}$ [^2] Other goals are possible such as reducing the source text complexity, removing specific information etc… Our choice stems from the practical usability of the task-oriented definition for evaluation purposes. Next, we formalize this evaluation metric based on the assumption of an unknown downstream task.
Let $c:\Omega^{*}\rightarrow\{1,\dots,m\}$ denote the target concept of interest which can be extracted from the initial texts $\mathbf{T}$ by applying $C\triangleq c(\mathbf{T})$; and similarly, let $\widehat{c}(\mathbf{S})$ denote the predicted concept from the summaries $\mathbf{S}$ according to the underlying model $p_{\theta}(\mathbf{s}|\mathbf{t})$. The evaluation of the summarizer’s quality with respect to the downstream task (unknown from the summarization model) can be assessed using the expected error rate, as illustrated in Figure 1. In other words, it involves the classifier, determining the average probability of the extracted concept $\widehat{c}(\mathbf{S})$ differing from the original concept $c(\mathbf{T})$ in the source text:
@@ 3.3 A task-oriented evaluation setting
where the expectation is taken with respect to the joint distribution of text and summary $(\mathbf{T},\mathbf{S})$ based on the source texts distribution and the distribution over the summaries induced by the summarizer. Interestingly, it is not difficult to check by data-processing inequality that (4): $I(C;\mathbf{S})\leqslant I(C;\mathbf{T})$ and thus, the output summaries $\mathbf{S}$ by the summarization model may not preserve all relevant information necessary to predict $C$ from $\mathbf{S}$, unless the summarization model is aware of the downstream task. However, identifying the set of relevant downstream tasks for various texts is challenging in practice. Consequently, (6) does not provide a satisfactory evaluation metric for summarization systems. This observation raises the central question studied in this work: How much information about any downstream task is preserved in the summaries?
4 A Task-Agnostic Quality Measure for Summarization
@@ 4.1 Theoretical Results
In this section, we rigorously motivate the evaluation of summarization systems by measuring the MI between texts and the resulting summaries. Let’s assume the existence of a random variable $\mathbf{T}$ representing source texts, which follows an unknown distribution inherent to the source text-domain. Given a stochastic summarization system $p_{\theta}(\mathbf{s}|\mathbf{t})$ as defined in expression (1), we denote $\mathbf{S}\sim p_{\theta}(\mathbf{s}|\mathbf{T})$ the random variable representing summaries generated by the summarization system for these sources texts $\mathbf{T}$.
Consider the task $c(\mathbf{T})$ that we intend to execute on the source texts, where $C=c(\mathbf{T})$ represents the random variable denoting the outcomes of this task on source texts. We assume that the same task can be performed on the summaries and let $\widehat{c}(\mathbf{S})$ denote the corresponding prediction. Intuitively, if we can accurately predict the outcome of $C$ from the summaries, then we will say that the summarization system is high-quality since it preserves the necessary information for performing the task. The next proposition frames our information-theoretic bounds on the performance of any arbitrary downstream task. In particular, it shows that the expected error rate $P_{e}(c,\widehat{c},\theta)$ as defined in (6) can be upper and lower bounded by the MI between the texts and the summaries.
@@
Let $C=c(\mathbf{T})$ denote the underlying concept variable and let $\widehat{c}(\mathbf{S})$ be the Bayes predictor of $C$ observing the output summaries $\mathbf{S}$, based on the underlying summarization model $p_{\theta}(\mathbf{s}|\mathbf{t})$. The expected error rate satisfies:
@@
The upper bound (7) relies on the fact that the predictor $\widehat{c}(\mathbf{S})$ is the optimal (Bayes) classifier for which the expected error rate admits a well-known expression. The lower bound (8) uses data-processing inequality which implies that $I(\mathbf{T};\mathbf{S})\geqslant I\big{(}c(\mathbf{T});\widehat{c}(\mathbf{S})\big{)}$ and the definition of the rate-distortion function evaluated in the loss $\ell_{01}(c,\widehat{c})$ noticing that its expectation yields the expected error rate $P_{e}(c,\widehat{c},\theta)$ as defined in (6). For further details, see Appendix A. ∎
@@
The bounds in Proposition 1 imply that the expected error rate of any task predicted by observing the summaries is lower- and upper-bounded by functions of the MI between the text and the summaries $I(\mathbf{T};\mathbf{S})$. More precisely, the upper bound (7) is a monotonically decreasing function of the MI while the lower bound (8) is a non-increasing function in the MI. The lower bound can be further simplified by evaluating the rate-distortion function, as shown in Appendix B. In other words, greater MI corresponds to improved the expected prediction performance on the summaries. Conversely, the expected error rate is bounded from above when MI is limited. Interestingly, the arguments regarding the bounds do not depend on the considered task, suggesting that the MI can be used as a task-agnostic metric to evaluate the quality of summaries.
In the next section, we propose a practical method to estimate MI. In Section 6, we empirically show that it correlates well with the performance of the downstream tasks with other human judgment-based metrics and compare it to standard metrics.
@@ 4.2 Estimating MI from samples
Mutual information estimator. We rely on the KNIFE estimator Pichler et al. (2022) to estimate the differential entropy of the embeddings and then mutual information. This estimator effectively relies on Gaussian Mixtures with $K$ modes to fit the density function and induces a soft-clustering for text generation evaluation Pillutla et al. (2021); Pimentel et al. (2023). We found experimentally that the number of modes $K$ did not impact the performance significantly. We report the results for $K=4$ (see Appendix D.1 for further details).
Embeddings. In order to obtain continuous representations of the source texts and the summaries, we mainly rely on the AnglE-embedders Li and Li (2023) as they are at the top of the MTEB leaderboard Muennighoff et al. (2023) with a rather small model. In addition, we experimented with different embedders from the sentence transformers library Reimers and Gurevych (2019), mainly the paraphrase and sentence similarity embeddings. We find that our method was robust in the choice of the embedder and thus reported the results for the AnglE-embedders.
@@ 5 Experimental Settings
While the bounds derived in Section 4.1 provide certain theoretical guarantees regarding the mutual information, their practical consequences must still be evaluated empirically. We present here two evaluation methods. First, we show that MI effectively predicts whether performing a downstream on the summaries would lead to the same outcome as if it were performed on the source text. This validates our task-oriented evaluation approach for summaries. Furthermore, we contrast our metric with metrics trained to emulate human judgments across various dimensions. Remarkably, we observe that even without any training, our metric closely aligns with human preferences.
Datasets. We select three summarization datasets for the English language: CNN/DailyMail See et al. (2017); Hermann et al. (2015), XSum Narayan et al. (2018) and MultiNews Fabbri et al. (2019) and perform all evaluations on all datasets. We provide additional experiments in French and Spanish using the MLSUM Scialom et al. (2020) and XLSUM Hasan et al. (2021) datasets. We report mixed results due to the lack of efficient multilingual embedders in Section E.2.
Models. We evaluate numerous summarizers from the HuggingFace hub, relying on different backbones, pretraining methods and finetuned on different datasets. We conduct our experiments on the PEGASUS suite of models Zhang et al. (2020a), on BERT large models Lewis et al. (2019) and on DistilBERT models Shleifer and Rush (2020). We also evaluated the models presented in the SEAHORSE benchmark Clark et al. (2023) through the generated summaries they proposed, which include MT5 models Xue et al. (2021), different variants of T5 models Raffel et al. (2023), and of the PaLM models Chowdhery et al. (2022). The reader is referred to Appendix C for models’ details.
@@
Human evaluation with SummEval. While the SummEval Fabbri et al. (2021) dataset provides only a very limited number of human judgments, it contains enough aligned source, and AI-generated summaries to evaluate the MI metric. We use these unannotated data to evaluate the MI and rank the model accordingly; then we compare this ranking with the human evaluation provided by the SummEval benchmark. We provide a comparison with previous metrics in Table 2. It is worth noting that the evaluation is not fair for our metric as other are evaluated only on the annotated data, while ours is evaluated on the unannotated data.